How many possible microstates are there for 1 mol of H2O (l) at room temperature? (Use S◦ = 188.8 J/K mol.)
Answer #1

For 1 mol

S = 188.8 J/K

then

S = k_\mathrm{B} \ln W \!   

which is Boltzmann equation and

1.38065 × 10−23 J/K.

W = number of microstates

then

188.8 = 1.38065 × 10−23 * lnW

W = exp((188.8)/(1.38*10^-23)) = exp(1.36811*10^25)

# microstates = exp(1.36811*10^25)